OpenStudy (anonymous):

Rhichard works at an ice cream shop. Regualr cones get two scoops of ice cream: large cones get three scoops. one hot saturday, Richard scooped 234 regular cones and 156 large cones. one scoop of ice cream is 3 ounces. a tub of ice cream weighs 10 pounds. how many tubs of ice cream did richard use to make the cones?

OpenStudy (anonymous):

total scoops = (234 x 2) + (156 x 3) Take this and multiply by 3oz per scoop = total oz. Divide that by 16 oz in a lb = total lbs. Divide that by 10 = number of tubs. You can do it, you should get 17.55

OpenStudy (anonymous):

answer options are a-936 tubs b-468 tubs c-175 1/2 tubs d- 11/20 tubs

OpenStudy (anonymous):

i dont know sry

OpenStudy (anonymous):

234 regular cones = 2(234) = 468 scoops. 1 scoop is 3 ounces, so 468 scoops = 1404 ounces. 156 large cones = 3(156) = 468 scoops. 1 scoop is 3 ounces, so 468 scoops = 1404 ounces So, you have a total of (1404 + 1404) = 2808 ounces. a tub of ice cream weighs 10 lbs. 1 lb = 16 ounces so 10 lb would be 160 ounces. each tub contains 160 ounces. Richard used 2808 ounces. Now divide 2808 by 160 = 17.55 I got the same thing as laurensimmons.....are you sure you wrote the right answer choices down ?

OpenStudy (anonymous):

@jim_thompson5910 ....what are we doing wrong ???

jimthompson5910 (jim_thompson5910):

another way to do it 1 tub = 10 pounds 1 tub = 10*16 oz 1 tub = 160 oz 1 tub = 160*(1 scoop/3 oz) 1 tub = 160/3 scoops 1 tub = 53.33333 scoops ---------------------------------- There are 234*2+156*3 = 936 scoops total So 936 scoops * (1 tub/ 53.33333 scoops) = 936/53.33333 = 17.550001 which is basically 17.55 because there was roundoff error

jimthompson5910 (jim_thompson5910):

so the answer is 17.55 you can represent this as 17 + 0.55 = 17 + 55/100 = 17 + 11/20 = 17 & 11/20

jimthompson5910 (jim_thompson5910):

which looks like choice D, assuming that you left off a 17 in that answer choice

OpenStudy (anonymous):

oh...so we were not wrong....you are good at math jim....thank you so much.

jimthompson5910 (jim_thompson5910):

you're welcome, you guys are great too

OpenStudy (anonymous):

either bryce forgot to put the 17 or the answer choices are not tubs, but scoops...then the answer would be 936 scoops ?

jimthompson5910 (jim_thompson5910):

well if it asks how many SCOOPS (and not tubs) are used, then yes, it's 936